m=-3;(-2,1)
Write the slope-intercept equation for the line with the given slope that contains the given point.

Found 3 solutions by stanbon, mananth, Bassam:
Answer by stanbon(75887)

m=-3;(-2,1)
Write the slope-intercept equation for the line with the given slope that contains the given point.
------------
Equation Form:
y = mx+b
Substitute for y,m,x, and solve for "b":
-----
1 = -3*-2 + b
1 = 6 + b
b = -5
----
Equation:
y = -3x-5
============
Cheers,
Stan H.
============

Answer by mananth(16949)

m=-3;(-2,1)

plug values in y = mx+b
1=-3(-2)+b
b=7
The equation is y= -3x+7

Answer by Bassam(2)

as you are given m=-3, slope/gradient=-3
Also m=(y-y1) / (x-x1)
therefore -3=y-(-2) / x-1 where y1=-2, x1=1, ie : (-2,1)
Cross multiply, -3(x-1)=y-(-2)
expand and simplify -3x+3=y+2
do -2 both side -3x+3-2=y+2-2
tidy up -3x+1=y
swap sides y=-3x+1
because the equation is in the form of y=mx+c
slope = co-efficient of x and intercept=c, constant
Again as you can see slope/gradient=-3 and intercept=1
I hope that answer is clear. Try doing it again and again. Also practise similar questions to understand more.